OpenStudy (anonymous):

(7 √2 -3 √3) (4√6+3√12)

OpenStudy (anonymous):

those symbols are supposed to be radicals... can anyone read that?

OpenStudy (displayerror):

Your problem: \[\left(7 \sqrt{2} - 3 \sqrt{3}\right) \times \left(4 \sqrt{6} + 3 \sqrt{12}\right)\] To solve this, you have to FOIL (acronym for First Outer Inner Last) Using a general expression \[\left(A + B\right) \times \left(C + D\right)\] FOIL is done like this: \[AC + AD + BC + BD \]

OpenStudy (anonymous):

so its broken down like..... 28 rad12 + 21 rad24 -12 rad 18 -9 rad 36?

OpenStudy (displayerror):

Yep, that's right! Now just combine the terms. Keep in mind that some radicals can be simplified further. As an example, \(\sqrt{18} = \sqrt{9 \times 2} = 3 \sqrt{2}\)

OpenStudy (anonymous):

thank you so much! that makes this so much easier!

OpenStudy (displayerror):

No problem! Happy studying!